# Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b

### v1

@@ 7.2 Prospects for atmospheric characterization
Here we provide a thorough examination of the simulated transmission spectra of planets utilizing PLanetary Atmospheric Transmission for Observer Noobs (PLATON; Zhang et al. 2019, 2020). Specifically, we employed PandExo (Batalha et al. 2017) to generate synthetic data, mimicking observations that would be made with the James Webb Space Telescope (JWST). Our simulation methodology encompasses various parameters such as planetary and stellar radii, effective temperatures, stellar metallicity, stellar gravity, transit duration, orbital period, and inclination. Through meticulous adjustment of these inputs, we created mock datasets to replicate transmission spectra. We ensured a minimum of $3$ transit observations, spanning between $3$-$4$ hours of observation time, with at least one-half of the time allocated for out-of-transit observations for each target to acquire the necessary simulated data. The resultant synthetic transmission spectra offer insights into the atmospheric properties of these planets. Through detailed simulation and analysis of these spectra, our goal is to contribute valuable information regarding the composition and characteristics of exoplanetary atmospheres. Figure 15 presents the synthetic spectra for TOI-5720 b, TOI-6008 b and TOI-6086 b.
The James Webb Space Telescope (JWST) is equipped with four scientific instruments (Gardner et al. 2006), delivering high-resolution and highly sensitive observations of celestial targets. The JWST NIRSpec (Rauscher et al. 2003) employs fixed slits (FS) spectroscopy, offering high sensitivity single-object spectroscopy across a wide wavelength range from 0.6 $\mu$m to 5.3 $\mu$m. The JWST NIRCam (Horner et al. 2002), in its wide-field slitless spectroscopy mode, utilizes grisms and filters covering wavelengths between 2.4 $\mu$m to 5.0 $\mu$m, achieving a resolving power of $1,600$ at 4 $\mu$m. Our study utilized the F322W2 and F [^4] W filters, spanning wavelengths from 2.7 $\mu$m to 4 $\mu$m and 4 $\mu$m to 5 $\mu$m, respectively. The JWST Near Infrared Imager and Slitless Spectrograph (NIRISS; (Doyon & Hutchings 2012)) provide slit-less spectroscopy across wavelengths ranging from 0.6 $\mu$m to 5.0 $\mu$m. The Mid-Infrared Instrument (MIRI; (Boccaletti et al. 2003)) extends our observational capabilities into the mid-infrared spectrum, covering spectral wavelengths from 4.9 $\mu$m to 28.8 $\mu$m.
In our investigation of three validated planets, we leverage various grism and filter combinations available for each instrument. These combinations act as instrumental tools facilitating a comprehensive simulation of the transmission spectra for each planet.
@@ 7.2.1 PLATON
PLATON (Zhang et al. 2019, 2020), a Python package, computes transmission spectra for exoplanets and derives atmospheric characteristics based on observed spectra. The software accommodates key atmospheric parameters, including temperature, metallicity, C/O ratio, cloud-top pressure, and scattering slope. Employing an equilibrium chemistry model computed with GGchem (Woitke et al. 2018), PLATON allows users to input metallicity and C/O ratio. From these inputs, PLATON calculates the abundances of $34$ chemical species in each atmospheric layer. GGchem computes these abundances across a grid of metallicity, C/O ratio, temperature, and pressure, assuming equilibrium chemistry with or without condensation. The parameter ranges are as follows: metallicity from $0.1$ to $1000$ times solar, C/O ratio from $0.2$ to $2$, temperature from $300$ to $3000$ K, and pressure from $10^{-4}$ to $10^{8}$ Pa. To compute a transmission spectrum, we adopt the assumption of an isothermal atmosphere with equilibrium chemistry. PLATON divides the atmosphere into 500 equally spaced layers in $\log{P}$ from $10^{-4}$ to $10^{8}$. The physical depth of each layer is determined by applying the hydrostatic equation. The opacity of each layer is subsequently calculated, taking into account factors such as gas absorption, collisional absorption, and Rayleigh scattering. The radiative transfer calculation is executed by considering a ray passing through the atmosphere with an impact parameter. The optical depth experienced by the ray is then computed, and the transit depth is derived by summing the cross-sectional areas of each layer, with weights assigned based on the amount of light permitted through. Subsequently, corrections for unocculted starspots are implemented, which are deemed crucial for active stars due to the distinct spectra of both hot, unspotted regions and colder starspots. The correction involves adjusting the transit depth based on factors such as the spot fraction, spot temperature, temperature of unspotted regions, and the spectrum of the stellar surface. Each layer’s composition is determined by considering equilibrium chemistry, which facilitates the calculation of the mean molecular weight. Fitted parameters include the mixing ratios of these molecules, the atmospheric mean molecular weight, the surface pressure, and the planet’s radius. The specified radius for the planet is taken at a reference pressure of $1$ bar, chosen for its proximity to the standard atmospheric pressure on Earth. Our model assumes a clear sky without clouds or haze.
@@ 7.2.2 PandExo
PandExo (Batalha et al. 2017) serves as both an online interface and an open-source Python package, employed in modeling instrumental noise for exoplanet transit time series spectroscopy modes with JWST (NIRISS, NIRCam, NIRSpec, and MIRI LRS) and HST (WFC3). Drawing on the foundations of the Space Telescope Science Institute’s (STScI) Exposure Time Calculator, Pandeia (Pontoppidan et al. 2016), PandExo exhibits excellent agreement, with deviations well within $10\%$ not only for HST WFC3 but also across the entire spectrum of JWST instruments. The input parameters encompass critical factors, including a stellar Spectral Energy Distribution (SED) model sourced from the Phoenix Stellar Atlas (Husser et al. 2013b), the apparent magnitude of the star, details about the planet’s spectrum (primary or secondary), transit duration, the ratio of time dedicated to in-transit versus out-of-transit observations, the number of observed transits, the saturation threshold expressed as a percentage of the full well capacity, and the user-defined noise floor. For the computation of the planet’s atmospheric transmission data, we utilized the built-in molecular lines grid in PandExo. Our approach involved selecting the planetary temperature closest to the computed temperature of each of the eight validated planets. We considered equilibrium chemistry with TiO and medium Rayleigh Scattering. It is essential to note that our modeling assumes conditions devoid of clouds and hazes, representing clear skies for these analyses.
Refer to captionRefer to caption
Figure 14: Feasibility of TOI-5720 b, TOI-6008 b and TOI-6086 b for transmission and emission spectroscopy studies. Transmission spectroscopy metric (TSM, top panel) and emission spectroscopy metric (ESM, bottom panel) as a function of planetary radius. TSM and ESM are computed using equations 1 and 4 from Kempton et al. (2018). Our planets are compared to the Earth-sized ($R_{p}<1.5$ R) known transiting exoplanets observed or scheduled on JWST around M-dwarfs ($T_{\rm eff}<3800$). The size of the points scale according to the stellar radius. The points are coloured according to the stellar effective temperature. The blue circles highlight our planets TOI-5720 b, TOI-6008 b and TOI-6086 b.
@@ 7.3 Results
In this section we discuss about the synthetic spectra we obtained using PandExo and best fit model we find using PLATON. In Table 6 we provided the details of exposure time and SNR values.
@@ 7.3.1 TOI-5720 b
Our model spectra using PLATON align well with the simulated data from PandExo. $\rm H_{2}O$ lines can be found at 1, 1.3, 1.8, and 2.8 $\mu$m using NIRSpec, NIRISS, and NIRCam. Observation using MIRI may also add a contribution to detecting subtle $\rm H_{2}O$ signatures at higher wavelengths. We also found $\rm CO_{2}$ signature at 4.3 $\mu$m, depending on the metallicity and C/O chosen for our study. Our fitting efforts yield a best-fit model with an estimated metallicity of the atmosphere of TOI-5720b to be 1.26 times Solar metallicity and C/O to be 1.01. We also find $\rm CH_{4}$ lines at 1.2, 1.4, 1.7, 2.3, and 3.3 $\mu$m with NIRSpec, NIRISS, and NIRCam (see top panel of Figure 15). There are also a possible presence of $\rm NH_{3}$ in small amount.
@@ 7.3.2 TOI-6008 b
Our model spectra using PLATON align well with the simulated data from PandExo. $\rm H_{2}O$ lines can be found at 0.8, 1.2, 1.4, 1.8, and 2.7 $\mu$m using NIRSpec, NIRISS, and NIRCam. Observation using MIRI may also add a contribution to detecting subtle $\rm H_{2}O$ signatures at higher wavelengths. We also found $\rm CO_{2}$ signature at 4.3 $\mu$m and $\rm CO$ at 4.8 $\mu$m, depending on the metallicity and C/O chosen for our study. Our fitting efforts yield a best-fit model with an estimated metallicity of the atmosphere of TOI-6008b to be 1.26 times Solar metallicity and C/O to be 0.5 (see middle panel of Figure 15).
@@ 7.3.3 TOI-6086 b
Our model spectra using PLATON align well with the simulated data from PandExo. $\rm H_{2}O$ lines can be found at about 0.8, 1.0, 1.3, 1.8, and 3 $\mu$m using NIRSpec, NIRISS, and NIRCam. Observation using MIRI may also add a contribution to detecting subtle $\rm H_{2}O$ signatures at higher wavelengths. We also found $\rm CO_{2}$ signature at 4.3 $\mu$m and $\rm CO$ at 4.8 $\mu$m, depending on the metallicity and C/O chosen for our study. Our fitting efforts yield a best-fit model with an estimated metallicity of the atmosphere of TOI-6086b to be 1.26 times Solar metallicity and C/O to be 0.3. There is also a possible small amount of $\rm NH_{3}$ present. We also find $\rm CH_{4}$ lines at 0.8, 1.2, 1.4, 1.7, 2.3, and 3.3 $\mu$m with NIRSpec, NIRISS, and NIRCam (see bottom panel of Figure 15).
Figure 15: Synthetic transmission spectrum in JWST for TOI-5720 b (top panel), TOI-6008 b (middle panel) and TOI-6086 b (bottom panel) generated with the help of PLATON and PandExo tools. Red, yellow, green and grey points are simulated data using JWST/MIRI, JWST/NIRSpec, JWST/NIRCam and JWST/NIRISS instruments, and the best-fitting model is superimposed in blue.
Instrument Mode a Exposure SNR
@@
a Details about different modes, resolving power and wavelength can be found on: https://jwst-docs.stsci.edu/
@@
c Size of sub array for single object slitless spectroscopy.
Table 6: Exposure times and SNR obtained for each synthetic specrum of JWST.
@@ 8 Software
This work made use of many publicly available packages & tools for which the authors are immensely grateful. In addition to the software and pipelines cited throughout the paper, we also used Matplotlib (Hunter 2007), Numpy (van der Walt et al. 2011), Astropy (Astropy Collaboration et al. 2013), SciPy (Jones et al. 2001), Wotan (Hippke et al. 2019), and Photutils (Bradley et al. 2020).
@@ 9 Acknowledgements
Appendix A Discovery TESS light curves for TOI-5720, TOI-6008 and TOI-6086
Figure 16: TESS  phase-folded photometric data of TOI-5720 b (blue), TOI-6008 b (purple) and TOI-6086 b (green), with best-fitting transit model is superimposed.
@@ Appendix A Discovery TESS light curves for TOI-5720, TOI-6008 and TOI-6086
We observed TOI-5720 on UT 2023 February 5 using the NN-EXPLORE Exoplanet Stellar Speckle Imager (NESSI; Scott et al. (2018)), a speckle imager employed at the WIYN 3.5 m telescope on Kitt Peak. NESSI was used to obtain simultaneous speckle imaging in two filters with central wavelengths $\lambda_{c}=562$ and 832 nm. The observation consisted of a set of 9 1000-frame 40 ms exposures. NESSI’s field-of-view was set by a $256\times 256$ pixel sub-array readout, which resulted in a $4.6\times 4.6$ arcsecond field. However, our speckle measurements were confined to an outer radius of 1.2 arcseconds from the target star. Speckle imaging of a point source standard star was also taken in conjunction to the observation of the TOI. The standard observation consisted of a single 1000-frame image set and was used to calibrate the underlying PSF.
These speckle data were reduced using a pipeline described by Howell et al. (2011). Among the pipeline products is a reconstructed image of the field around TOI-5720 in each filter. The reconstructed image was used to measure a contrast curve setting the detection limits on any point sources in close proximity to TOI-5720. The contrast level was obtained by measuring fluctuations in the noise-like background level as a function of separation from the target star. No companion sources were detected for TOI-5720 in the NESSI data (see Figure 17).
Figure 17: High resolution imaging of TOI-5720 obtained with the WIYN-3.5m/NESSI on UT 2023 Feb 5 in two filters with central wavelengths $\lambda_{c}=562$ and 832 nm. No close companion sources were detected near the target.
@@ Appendix C High-Resolution imaging: Robo AO
We used additional sub-arcsecond imaging taken with Robo-AO, an autonomous laser-guided adaptive optics (AO) system (Baranec et al. 2014). The Robo-AO M-dwarf multiplicity catalog contains AO imaging of TOI-5720 and TOI-6008, taken on 2016 February 18 and 2016 June 18, respectively, with the 2.1m Kitt Peak telescope (Lamman et al. 2020). The images were taken with an Andor iXon DU-888 camera in the $i^{\prime}$ filter with an exposure time of 90s. Median seeing at the telescope was 1.44” which resulted in a FWHM of around 0.12”. Observations were then processed with automatic pipelines, including one optimized for either high or low signal-to-noise images and a ”high-contrast” imaging pipeline which creates a locally-optimized PSF-subtracted image around the target using reference starts observed the same night (Jensen-Clem et al. 2018). Contrast curves of the observations are shown in Figure 18; there is no indication of any additional close stellar companions were detected above these contrast curves.
@@
Appendix D Posterior probability distribution for the TOI-5720, TOI-6008 and TOI-6086 systems
Figure 19: The PDF (Posterior probability distribution) for the TOI-5720 system parameters derived from our global MCMC analysis. The median value for each parameter is represented by the vertical dashed lines.
Figure 20: The PDF for the TOI-6008 system.
Figure 21: The PDF for the TOI-6086 system.

### v2

@@ 7.2 Prospects for atmospheric characterization
We provide a thorough examination of the simulated transmission spectra of planets using PLanetary Atmospheric Transmission for Observer Noobs (PLATON; Zhang et al. 2019, 2020). Specifically, we employed PandExo (Batalha et al. 2017) to generate synthetic data, mimicking observations that would be made with the James Webb Space Telescope (JWST). Our simulation method encompasses various parameters such as planetary and stellar radii, effective temperatures, stellar metallicity, stellar gravity, transit duration, orbital period, and inclination. Through meticulous adjustment of these inputs, we created mock datasets to replicate transmission spectra. We ensured a minimum of three transit observations, spanning between $3$-$4$ hours of observation time, with at least one-half of the time allocated for out-of-transit observations for each target to acquire the necessary simulated data. The resultant synthetic transmission spectra offer insights into the atmospheric properties of these planets. Through detailed simulation and analysis of these spectra, our goal is to contribute valuable information regarding the composition and characteristics of exoplanetary atmospheres. Figure 14 presents the synthetic spectra for TOI-5720 b, TOI-6008 b, and TOI-6086 b.
The James Webb Space Telescope (JWST) is equipped with four scientific instruments (Gardner et al. 2006), delivering high-resolution and highly sensitive observations of celestial targets. The JWST NIRSpec (Rauscher et al. 2003) employs fixed-slit (FS) spectroscopy, offering high-sensitivity single-object spectroscopy for a wide wavelength range from 0.6 $\mu$m to 5.3 $\mu$m. The JWST NIRCam (Horner et al. 2002), in its wide-field slitless spectroscopy mode, uses grisms and filters covering wavelengths between 2.4 $\mu$m to 5.0 $\mu$m, achieving a resolving power of $1,600$ at 4 $\mu$m. Our study used the F322W2 and F [^4] W filters, spanning wavelengths from 2.7 $\mu$m to 4 $\mu$m and 4 $\mu$m to 5 $\mu$m, respectively. The JWST Near Infrared Imager and Slitless Spectrograph (NIRISS; (Doyon & Hutchings 2012)) provides slit-less spectroscopy across wavelengths ranging from 0.6 $\mu$m to 5.0 $\mu$m. The Mid-Infrared Instrument (MIRI; (Boccaletti et al. 2003)) extends our observational capabilities into the mid-infrared spectrum, covering spectral wavelengths from 4.9 $\mu$m to 28.8 $\mu$m.
In our investigation of the three validated planets, we used various grism and filter combinations available for each instrument. These combinations act as instrumental tools facilitating a comprehensive simulation of the transmission spectra for each planet.
@@ 7.2.1 PLATON
PLATON (Zhang et al. 2019, 2020) is a Python package that computes transmission spectra for exoplanets and derives atmospheric characteristics based on observed spectra. The software accommodates key atmospheric parameters, including temperature, metallicity, C/O ratio, cloud-top pressure, and scattering slope. Employing an equilibrium chemistry model computed with GGchem (Woitke et al. 2018), PLATON allows users to input the metallicity and C/O ratio. From these inputs, PLATON calculates the abundances of $34$ chemical species in each atmospheric layer. GGchem computes these abundances across a grid of metallicity, C/O ratio, temperature, and pressure, assuming equilibrium chemistry with or without condensation. The parameter ranges are as follows: a metallicity from $0.1$ to $1000$ times solar, a C/O ratio from $0.2$ to $2$, a temperature from $300$ to $3000$ K, and a pressure from $10^{-4}$ to $10^{8}$ Pa. To compute a transmission spectrum, we adopted the assumption of an isothermal atmosphere with equilibrium chemistry. PLATON divided the atmosphere into 500 equally spaced layers in $\log{P}$ from $10^{-4}$ to $10^{8}$. The physical depth of each layer was determined by applying the hydrostatic equation. The opacity of each layer was subsequently calculated, taking into account factors such as gas absorption, collisional absorption, and Rayleigh scattering. The radiative transfer calculation was executed by considering a ray passing through the atmosphere with an impact parameter. The optical depth experienced by the ray was then computed, and the transit depth was derived by summing the cross-sectional areas of each layer, with weights assigned based on the amount of light that was permitted to pass through. Subsequently, corrections for unocculted starspots were implemented, which were deemed crucial for active stars due to the distinct spectra of hot, unspotted regions and colder starspots. The correction involved adjusting the transit depth based on factors such as the spot fraction, the spot temperature, the temperature of unspotted regions, and the spectrum of the stellar surface. The composition of each layer was determined by considering equilibrium chemistry, which facilitated the calculation of the mean molecular weight. The fitted parameters included the mixing ratios of these molecules, the atmospheric mean molecular weight, the surface pressure, and the planet radius. The specified radius for the planet was taken at a reference pressure of $1$ bar, which was chosen for its proximity to the standard atmospheric pressure on Earth. Our model assumed a clear sky without clouds or haze.
@@ 7.2.2 PandExo
PandExo (Batalha et al. 2017) serves as both an online interface and an open-source Python package, which we employed to model instrumental noise for exoplanet transit time-series spectroscopy modes with JWST (NIRISS, NIRCam, NIRSpec, and MIRI LRS) and HST (WFC3). Drawing on the foundations of the Space Telescope Science Institute (STScI) Exposure Time Calculator, Pandeia (Pontoppidan et al. 2016), PandExo exhibits excellent agreement, with deviations well within $10\%$ not only for HST WFC3, but also across the entire spectrum of the JWST instruments. The input parameters encompass critical factors, including a stellar SED model sourced from the Phoenix stellar atlas (Husser et al. 2013b), the apparent magnitude of the star, details of the planet spectrum (primary or secondary), the transit duration, the ratio of time dedicated to in-transit versus out-of-transit observations, the number of observed transits, the saturation threshold expressed as a percentage of the full-well capacity, and the user-defined noise floor. To compute the atmospheric transmission data of the planet, we used the built-in molecular line grid in PandExo. Our approach involved selecting the planetary temperature closest to the computed temperature of each of the eight validated planets. We considered equilibrium chemistry with TiO and medium Rayleigh scattering. It is essential to note that our modeling assumed conditions devoid of clouds and hazes, representing clear skies for these analyses.
Figure 13: Feasibility of TOI-5720 b, TOI-6008 b, and TOI-6086 b for transmission and emission spectroscopy studies. Transmission spectroscopy metric (TSM, top panel) and emission spectroscopy metric (ESM, bottom panel) as a function of planetary radius. Our planets are compared to the Earth-sized ($R_{p}<1.5$ R) known transiting exoplanets observed or scheduled on JWST around M dwarfs ($T_{\rm eff}<3800$). The size of the points scales with the stellar radius. The points are colored according to the effective stellar temperature. The blue circles highlight our planets TOI-5720 b, TOI-6008 b, and TOI-6086 b.
@@ 7.3 Results
We discuss about the synthetic spectra we obtained using PandExo and the best-fit model we found using PLATON. In Table 6 we provide the details of exposure time and S/N values.
@@ 7.3.1 TOI-5720 b
Our model spectra using PLATON align well with the simulated data from PandExo. $\rm H_{2}O$ lines can be found at 1, 1.3, 1.8, and 2.8 $\mu$m using NIRSpec, NIRISS, and NIRCam. Observation using MIRI may also add a contribution to detecting subtle $\rm H_{2}O$ signatures at higher wavelengths. We also found a $\rm CO_{2}$ signature at 4.3 $\mu$m, depending on the metallicity and C/O chosen for our study. Our fitting efforts yield a best-fit model with an estimated metallicity of the atmosphere of TOI-5720b of 1.26 times solar metallicity and C/O of 1.01. We also find $\rm CH_{4}$ lines at 1.2, 1.4, 1.7, 2.3, and 3.3 $\mu$m with NIRSpec, NIRISS, and NIRCam (see top panel of Figure 14). $\rm NH_{3}$ may also be present in a small amount.
@@ 7.3.2 TOI-6008 b
Our model spectra using PLATON align well with the simulated data from PandExo. $\rm H_{2}O$ lines can be found at 0.8, 1.2, 1.4, 1.8, and 2.7 $\mu$m using NIRSpec, NIRISS, and NIRCam. Observation using MIRI may also add a contribution to detecting subtle $\rm H_{2}O$ signatures at higher wavelengths. We also found $\rm CO_{2}$ signatures at 4.3 $\mu$m and $\rm CO$ at 4.8 $\mu$m, depending on the metallicity and C/O chosen for our study. Our fitting efforts yield a best-fit model with an estimated metallicity of the atmosphere of TOI-6008b of 1.26 times solar metallicity and C/O of 0.5 (see middle panel of Figure 14).
@@ 7.3.3 TOI-6086 b
Our model spectra using PLATON align well with the simulated data from PandExo. $\rm H_{2}O$ lines can be found at about 0.8, 1.0, 1.3, 1.8, and 3 $\mu$m using NIRSpec, NIRISS, and NIRCam. Observation using MIRI may also add a contribution to detecting subtle $\rm H_{2}O$ signatures at higher wavelengths. We also found $\rm CO_{2}$ signatures at 4.3 $\mu$m and $\rm CO$ at 4.8 $\mu$m, depending on the metallicity and C/O chosen for our study. Our fitting efforts yield a best-fit model with an estimated metallicity of the atmosphere of TOI-6086b of 1.26 times solar metallicity and C/O of 0.3. $\rm NH_{3}$ may also be present in a small amount of present. We also find $\rm CH_{4}$ lines at 0.8, 1.2, 1.4, 1.7, 2.3, and 3.3 $\mu$m with NIRSpec, NIRISS, and NIRCam (see bottom panel of Figure 14).
Figure 14: Synthetic transmission spectrum in JWST for TOI-5720 b (top panel), TOI-6008 b (middle panel) and TOI-6086 b (bottom panel) generated with the help of the PLATON and PandExo tools. The red, yellow, green, and grey points are simulated data using Jthe WST/MIRI, JWST/NIRSpec, JWST/NIRCam, and JWST/NIRISS instruments, and the best-fitting model is superimposed in blue.
Instrument Mode a Exposure S/N
@@
a Details of the different modes, resolving power and wavelength can be found at: https://jwst-docs.stsci.edu/
@@
c Size of the subarray for single-object slitless spectroscopy.
Table 6: Exposure times and S/N obtained for each synthetic specrum of JWST.
@@ 8 Software
This work made use of many publicly available packages and tools for which the authors are immensely grateful. In addition to the software and pipelines cited throughout the paper, we also used Matplotlib (Hunter 2007), Numpy (van der Walt et al. 2011), Astropy (Astropy Collaboration et al. 2013), SciPy (Jones et al. 2001), Wotan (Hippke et al. 2019), and Photutils (Bradley et al. 2020).
@@ 9 Acknowledgements
Appendix A Discovery TESS light curves for TOI-5720, TOI-6008, and TOI-6086
Figure 15: TESS phase-folded photometric data of TOI-5720 b (blue), TOI-6008 b (purple), and TOI-6086 b (green). The best-fitting transit model is superimposed.
@@ Appendix A Discovery TESS light curves for TOI-5720, TOI-6008, and TOI-6086
We observed TOI-5720 on UT 2023 February 5 using the NN-EXPLORE Exoplanet Stellar Speckle Imager (NESSI; Scott et al. (2018)), a speckle imager employed at the WIYN 3.5 m telescope on Kitt Peak. NESSI was used to obtain simultaneous speckle imaging in two filters with central wavelengths of $\lambda_{c}=562$ and 832 nm. The observation consisted of a set of nine 1000-frame 40 ms exposures. The NESSI field of view was set by a $256\times 256$ pixel subarray readout, which resulted in a field of $4.6\times 4.6$ arcsecond. However, our speckle measurements were confined to an outer radius of 1.2 arcseconds from the target star. Speckle imaging of a point source standard star was also taken in conjunction with the observation of the TOI. The standard observation consisted of a single 1000-frame image set and was used to calibrate the underlying PSF.
These speckle data were reduced using a pipeline described by Howell et al. (2011). The pipeline products include a reconstructed image of the field around TOI-5720 in each filter. The reconstructed image was used to measure a contrast curve setting the detection limits on any point source in close proximity to TOI-5720. The contrast level was obtained by measuring fluctuations in the noise-like background level as a function of separation from the target star. No companion sources were detected for TOI-5720 in the NESSI data (see Figure 16).
Figure 16: High-resolution imaging of TOI-5720 obtained with the WIYN-3.5m/NESSI on UT 2023 Feb 5 in two filters with central wavelengths $\lambda_{c}=562$ and 832 nm. No close companion sources were detected near the target.
@@ Appendix C High-resolution imaging: Robo AO
We used additional subarcsecond imaging taken with Robo-AO, an autonomous laser-guided adaptive optics (AO) system (Baranec et al. 2014). The Robo-AO M-dwarf multiplicity catalog contains AO imaging of TOI-5720 and TOI-6008, taken on 2016 February 18 and 2016 June 18, respectively, with the 2.1m Kitt Peak telescope (Lamman et al. 2020). The images were taken with an Andor iXon DU-888 camera in the $i^{\prime}$ filter with an exposure time of 90s. Thee median seeing at the telescope was 1.44”, which resulted in an FWHM of about 0.12”. Observations were then processed with automatic pipelines, including one pipeline that was optimized for images with either high or low signal-to-noise ratios and a high-contrast imaging pipeline that creates a locally optimized PSF-subtracted image around the target using reference stars observed on the same night (Jensen-Clem et al. 2018). The contrast curves of the observations are shown in Figure 17; there is no indication that any additional close stellar companions were detected above these contrast curves.
@@
Appendix D Posterior probability distribution for the systems TOI-5720, TOI-6008, and TOI-6086.
Figure 18: Posterior probability distribution for the TOI-5720 system parameters derived from our global MCMC analysis. The median value for each parameter is represented by the vertical dashed lines.
Figure 19: Posterior probability distribution for the TOI-6008 system.
Figure 20: Posterior probability distribution for the TOI-6086 system.